# WorldLLM: Improving LLMs’ world modeling using curiosity-driven theory-making

### v1

@@ B.2 Scientist
You are in an environment that contains multiple objects. It can contain water, plant seeds(carrot, porato, beet, berry and pea seeds), small herbivores(pig, cow and ship) and large herbivores(elephant, giraffe, rhinoceros). You can move objects, like water, plants or herbivores and release them on another object to make them interact and transform into a new object. Your previous experiences were:
@@ B.3.2 RL agents
Concerning the reward signals, we provide in Algorithm 2 the details on how RL-ALPEXP compute its ALP reward. RL-ALPEXP introduces an additional hyperparameter $\alpha$ which regulates the reward’s smoothing of the reward. For each transition type, $\alpha$ corresponds to the weight of the previous rewards obtained for this transition type. Its role is highlighted in our study from Appendix D.4.1 but $\alpha$ can be interpreted as an adjustment factor determining the extent to which rewards from previous steps are retained. A value of 00 for $\alpha$ implies that only the reward from the most recent step is considered, which corresponds to RL-ALP, while a value of $1$ indicates that the rewards are accumulated from the beginning. After a manual hyperparameter search (see Appendix D.4.2), we chose $\alpha=0.9$ across all transition types.
@@
2:Initialize $(m_{i})_{i\in\mathcal{I}}\leftarrow 0$triangleright Initialize moving average for each type
@@
7:         Compute reward $(r_{alp})_{\tau}$triangleright Compute ALP reward using equation 2
@@ Appendix D Additional Results
In this section, we first provide the complete evolution of log-likehoods on our test set from Section 3.3. Then, we provide additional studies and experiments to improve one’s comprehension of WorldLLM’s behavior. These investigations notably explore the reasons behind the failure of RL-ALP and its differences with RL-ALPEXP.
@@ D.2.2 Embeddings of the hypotheses
In order to compare the hypotheses generated by our Scientist, we used a sentence transformer, all-mpnet-base-v2333https://huggingface.co/sentence-transformers/all-mpnet-base-v2, to project hypotheses in a latent space. We then used T-SNE(Maaten & Hinton, 2008) to obtain a 2D representation of the embeddings. As the number of generated hypotheses is high, we plotted only 1/10 of them for each Experimenter and seed. We show these embeddings in Figure 13.
@@ D.4 A deeper study RL-ALP
In this section, we provide a deeper analysis of RL-ALP’s failure. While expected to be more robust than RL-LogP (e.g. to the "noisy TV" problem), our results showed RL-ALP was unable to collect other transitions than Standing, leading to very poor hypotheses. We begin this section by studying how directly using the per-transition ALP as reward for an RL policy is challenging. Then, show how RL-ALPEXP differs and helps stabilizing training.
@@ D.4.1 Failure of RL-ALP as the reward
To get a better grasp of the RL policy’s behavior when using RL-ALP, we plot the evolution of rewards obtained by the policy in Figure 15. We observe that most rewards equal 00. Higher rewards are only obtained in the first data collections when Standing transitions are not well predicted by the Statistician. After gathering such transitions, the policy eventually converges to a random one. This absence of rewards can be explained by our reward definition from Eq. 2.

### v2

@@ B.2 Scientist
You are in an environment that contains multiple objects. It can contain water, plant seeds(carrot, potato, beet, berry and pea seeds), small herbivores(pig, cow and ship) and large herbivores(elephant, giraffe, rhinoceros). You can move objects, like water, plants or herbivores and release them on another object to make them interact and transform into a new object. Your previous experiences were:
@@ B.3.2 RL agents
Concerning the reward signals, we provide in Algorithm 2 the details on how RL-ALPEXP compute its ALP reward. RL-ALPEXP introduces an additional hyperparameter $\alpha$ which regulates the reward’s smoothing of the reward. For each transition type, $\alpha$ corresponds to the weight of the previous rewards obtained for this transition type. Its role is highlighted in our study from Appendix D.4.1 but $\alpha$ can be interpreted as an adjustment factor determining the extent to which rewards from previous steps are retained. A value of 0 for $\alpha$ implies that only the reward from the most recent step is considered, which corresponds to RL-ALP, while a value of $1$ indicates that the rewards are accumulated from the beginning. After a manual hyperparameter search (see Appendix D.4.2), we chose $\alpha=0.9$ across all transition types.
@@
2:Initialize $(m_{i})_{i\in\mathcal{I}}\leftarrow 0$ $\triangleright$ Initialize moving average for each type
@@
7:   Compute reward $(r_{alp})_{\tau}$
@@ Appendix D Additional Results
In this section, we first provide the complete evolution of log-likelihoods on our test set from Section 3.3. Then, we provide additional studies and experiments to improve one’s comprehension of WorldLLM’s behavior. These investigations notably explore the reasons behind the failure of RL-ALP and its differences with RL-ALPEXP.
@@ D.2.2 Embeddings of the hypotheses
In order to compare the hypotheses generated by our Scientist, we used a sentence transformer, all-mpnet-base-v2333https://huggingface.co/sentence-transformers/all-mpnet-base-v2, to project hypotheses in a latent space. We then used T-SNE(maaten_visualizing_2008) to obtain a 2D representation of the embeddings. As the number of generated hypotheses is high, we plotted only 1/10 of them for each Experimenter and seed. We show these embeddings in Figure 13.
@@ D.4 A deeper study RL-ALP
In this section, we provide a deeper analysis of RL-ALP’s failure. While expected to be more robust than RL-LogP (e.g. to the "noisy TV" problem), our results showed RL-ALP was unable to collect other transitions than Standing, leading to very poor hypotheses. We begin this section by studying how directly using the per-transition ALP as reward for an RL policy is challenging. Then, show how RL-ALPEXP differs and helps stabilize training.
@@ D.4.1 Failure of RL-ALP as the reward
To get a better grasp of the RL policy’s behavior when using RL-ALP, we plot the evolution of rewards obtained by the policy in Figure 15. We observe that most rewards equal 0. Higher rewards are only obtained in the first data collections when Standing transitions are not well predicted by the Statistician. After gathering such transitions, the policy eventually converges to a random one. This absence of rewards can be explained by our reward definition from Eq. 2.